Text can be centred on the page by putting it within the center environment, and it will appear flush against the left or right margins if it is placed within the flushleft or flushright environments. Notice the spelling of center—unfortunately L^{A}TEX doesn’t understand the English spelling.

Text within these environments will be formatted in the normal way, in particular the ends of the lines that you type are just regarded as spaces. To indicate a “newline” you need to type the \\ command. For example:

     \begin{center}
     one
     two
     three \\
     four \\
     five
     \end{center}

one two three four five